# Sub-nanometer 3D morphometric precision of polarisation-resolved wide-field optical extinction microscopy determines the roundness of individual gold nanospheres

### v1

@@ IV.2 Morphometric analysis
In Fig. 4c, we examine the uncertainty in the retrieved ARs and $D_{V}$ due to measurement noise for four selected NPs, indicated by the colored stars in (a–b). The analysis was done for all three permittivity datasets with $g$ minimizing $\bar{S}$. We simulate the effect of the measurement noise by adding a realisation of Gaussian random noise with a standard deviation of $\hat{\sigma}_{\Lambda}(\gamma_{\mathrm{P}})$ (see inset of Fig. 2a) to the measured $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$, and then repeat the morphometric analysis. Each set of scatter data in each of the four panels represents 1000 noise realizations. The results for OL excluding $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ typically show the densest clusters, but often multiple displaced by up to $0.15$ in the ARs. Including $\sigma_{\Lambda}(\gamma_{\mathrm{R}})$ consistently retrieves dense clusters with the small separation. For the ARs, the width of the retrieved distributions is typically below $5\%$. For near-spherical NPs ARs retrieval is complicated by dominating quadratic parameter dependencies close to symmetry points of the the multi-dimensional parameter space.[43] More asymmetric NPs instead yield more precise retrieval, as exemplified by the blue-starred NP in (c), which presents one of the strongest observed non-sphericity in the sample. In this case, precision is better than $3\%$, corresponding to 0.84 nm length along the NP axes. For further insights into this effect, see Sec. LABEL:SM-sec:optnoise. Variations in $n_{\mathrm{m}}$ result in systematic changes in $c/b$ of about twice the index change, as discussed in Sec. LABEL:SM-sec:fitindex. For a realistic index accuracy of 0.01, the correspoding $c/b$ systematic error is 2%. Remarkably, the retrieved $D_{V}$ shows $0.25$ nm precision.
@@
The systematic deviation of about $2\,\mathrm{nm}$ between the mean of the retrieved $D_{V}$ and that found by TEM might be due to several factors. A known systematic error is the quasi-static approximation used in the model. Considering the wavelength in vacuum of the LSPR for a spherical gold NP of about $\lambda_{0}=530\,\mathrm{nm}$, the wavelength in the medium of $n_{\mathrm{m}}=1.52$ is $\lambda_{0}/n_{\mathrm{m}}\approx 350\,\mathrm{nm}$, which is about 12 times the nominal diameter of the NPs. This suggests that taking into account retardation, which leads to radiative damping and red-shift of the LSPR, might be relevant to achieve accuracies below 10% for the investigated NPs. Indeed, the extinction spectrum of a gold NP with the TEM-measured average diameter D=27.94D=27.94\,nm, calculated including retardation using Mie theory (see inset of Fig. 2a, dotted black line), shows a reduced cross-section at the LSPR and lower wavelengths. The vertical lines in Fig. 2a show the corresponding weighted cross-sections (solid for Rayleigh-Gans, dotted for Mie) for all colour channels. We see that particularly in the blue channels ($\color[rgb]{0,0,1}450\,\mathrm{nm}$}, $\color[rgb]{0,0.92,0.92}500\,\mathrm{nm}$}), the Mie results align with the measured extinction distributions, while the Rayleigh-Gans results are too large. From these observations we can expect that including retardation would reduce the systematic deviation of the mean diameter to below 1 nm, comparable to the precision. This indicates that the ON experimental systematic errors are small, and that the Rayleigh-Gans model presently used is limiting the accuracy for the investigated NPs.
@@ V Conclusions
We have presented an enhanced ON method with optical cross-sections measured using six spectral channels using both linear and radial polarisation to probe both in-plane and out-of-plane NP response. Measuring $30\,\mathrm{nm}$ ultra-uniform gold nanospheres, we demonstrated significant improvement in the precision and accuracy of the method. Precision in the size and shape has been improved to $\pm 0.25\,\mathrm{nm}$ and $5\%$, respectively. The accuracy is within 10% in shape and 8% in size using TEM measurements as reference. We examined the effect of the permittivity dataset on the morphometry retrieval and found that Olmon et al.[47] with surface damping factor $g=1.8$ yields the best results. We showed that the accuracy in size is limited by the Rayleigh-Gans model employed. Including retardation, as shown for spherical particles using Mie theory, allows to improve the accuracy to sub-nm. Future development could see the use of numerical models, capable of treating particles outside of the limitations analytical models. Another approximation which could be limiting is the use of a local permittivity to describe the nanoparticle optical response, which is known to break down on the sub-nm scale in general. Again, more complex theory models[56] can be used to improve on this.
The precision of the method is so high that small systematic effects in the experimental setup, such as the angular resolved transmission and the angular intensity distributions, are relevant to achieve an accuracy comparable of better than the precision. These can be corrected by further characterisation of the instrumentation.
Overall, this study shows the potential of the ON not only as as a complement to TEM, but also as a standalone methodology for accurate NP morphometric studies. It also offers an interesting approach to determining the permittivity of materials on the nanoscale. Specifically, it could be used to uncover the effects of non-local permittivities to describe nanoparticle optical response.

### v2

@@ IV.2 Morphometric analysis
In Fig. 4c, we examine the uncertainty in the retrieved ARs and $D_{V}$ due to measurement noise for four selected NPs, indicated by the colored stars in Fig. 4a. The analysis was done for all three permittivity datasets with $g$ minimizing $\bar{S}$. We simulate the effect of the measurement noise by adding a realisation of Gaussian random noise with a standard deviation of $\hat{\sigma}_{\Lambda}(\gamma_{\mathrm{P}})$ (see inset of Fig. 2a) to the measured $\sigma_{\Lambda}(\gamma_{\mathrm{P}})$, and then repeat the morphometric analysis. Each set of scatter data in each of the four panels represents 1000 noise realizations. The ARs often exhibit bimodal distributions, with the greatest separation between modes apparent for MP and JC, while the width of the individual modes is typically below $5\%$. Notably, the modes often are close to $b=a$ or $c=b$, allowing to achieve zero in-plane asymmetry (which is defined with high precision by the linear polarisation data) for appropriate NP orientations. Using synthetic $\sigma_{\mathrm{ext}}$ data generated using the Rayleigh-Gans model and the specified experimental corrections, such multimodal results are not observed,[43] suggesting that they are caused by systematic measurement errors. Notably, the sample distributions in Fig. 4a for each permittivity mirrors the spread of the noise distributions for the rounder NPs in Fig. 4c for each permittivity, which points to a homogeneity of the sample, as well as this systematic effect. We will return to this topic again later in the text. Generally, for near-spherical NPs, AR retrieval is less precise due to the quadratic parameter dependencies close to symmetry points of the the multi-dimensional parameter space.[43] Significantly asymmetric NPs instead yield more precise retrieval, as exemplified by the blue-starred NP. In this case, precision is better than $3\%$, corresponding to 0.84 nm length along the NP axes. For further insights into this effect, see Sec. C. Variations in $n_{\mathrm{m}}$ result in systematic changes in $c/b$ of about twice the index change, as discussed in Sec. F. For a realistic refractive index accuracy of 0.01, the corresponding $c/b$ systematic error is 2%. Remarkably, the retrieved $D_{V}$ shows a precision of $(0.04-0.1)\,\mathrm{nm}$ standard deviation for the selected NPs.
@@
The systematic deviation of about $2\,\mathrm{nm}$ between the mean of the retrieved $D_{V}$ and that found by TEM might be due to several factors. A known systematic error is the quasi-static approximation used in the model. Considering the wavelength in vacuum of the LSPR for a spherical gold NP of about $\lambda_{0}=530\,\mathrm{nm}$, the wavelength in the medium of $n_{\mathrm{m}}=1.52$ is $\lambda_{0}/n_{\mathrm{m}}\approx 350\,\mathrm{nm}$, which is about 12 times the nominal diameter of the NPs. This suggests that taking into account retardation, which leads to radiative damping and red-shift of the LSPR, might be relevant to achieve accuracies below 10% for the investigated NPs. Indeed, the extinction spectrum of a gold NP with the TEM-measured average diameter $D_{V}=27.94\,$nm, calculated including retardation using Mie theory (see inset of Fig. 2a, dotted black line), shows an about 4% increase and 5 nm red-shift of the LSPR peak, see also Sec. G. The vertical lines in Fig. 2a show the corresponding weighted cross-sections (solid for Rayleigh-Gans, dotted for Mie) for all colour channels. We would expect that the redshift would result in rounder retrieved NPs, while the increase would result in an about 1% smaller $D_{V}$ considering the approximate scaling of the cross-section with the volume.
Figure 5: Similar to Fig. 4, but for OL and MP using $\epsilon_{\mathrm{NP}}^{s}(\lambda)=\epsilon_{\mathrm{NP}}(\lambda+\delta\lambda)$ with $\delta\lambda=5\,\mathrm{nm}$ and the $0.97\times$ correction for $\sigma_{\mathrm{ext}}(\gamma_{\mathrm{R}})$. (a) Retrieved ARs shown as filled symbols (MP triangles, OL circles) with a color indicating the fitted diameter $D_{V}$. (b) MAR (empty symbols) and PAR (filled symbols) versus $D_{\mathrm{p}}$, with the histogram showing the number distribution of $D_{\mathrm{p}}$, with OL shown as magenta circles / bars and MP as cyan triangles / bars. The black filled squares indicate TEM measurements PAR versus $D_{\mathrm{p}}$. Inset: pitch ($\theta$) and roll ($\phi$) angles showing the particle orientation. (c) Simulated effect of the measurement noise on the retrieved ARs for individual NPs indicated by the colored stars in (a). Results for 1000 realisations of Gaussian noise with $\hat{\sigma}_{\mathrm{ext}}$ standard deviation added to the measured extinction data are shown. Inset: $D_{V}$ histograms, providing a colour scale for the symbols.
To investigate this further while keeping the Rayleigh-Gans model, we approximate the effects of retardation by spectrally shifting the permittivity with respect to $\lambda$ to obtain $\epsilon_{\mathrm{NP}}^{s}(\lambda)=\epsilon_{\mathrm{NP}}(\lambda+\delta\lambda)$, with $\delta\lambda=-5\,\mathrm{nm}$, and repeating the ON analysis using $\epsilon_{\mathrm{NP}}^{s}$ and the $0.97\times$ correction for $\sigma_{\mathrm{ext}}(\gamma_{\mathrm{R}})$. The resulting dependence of $\bar{S}$ and the median MAR on $g$ is given as dotted lines in Fig. 3a for OL and MP, showing that the $g$ at which $\bar{S}$ reaches a minimum is rather unaffected, while the median MAR increases. The corresponding ON results are given in Fig. 5 for OL and MP. The retrieved shapes are rounder for both datasets, with OL showing clearly round shape retrieval with median MAR of $0.954$, compared to $0.892$ for MP. Meanwhile, the median diameter for MP is $M\{D_{V}\}=26.72\,\mathrm{nm}$, larger than $25.77\,\mathrm{nm}$ for OL. For OL, nearly all of the asymmetry is in $b/a$  corresponding to a prolate particle, and most orientations show $\theta\approx-\pi/2$, corresponding to a prolate NP standing on end, i.e. its long axis is normal to the surface. This suggests a systematic effect, where a negligible in-plane polarisation-dependence combines with a constrain by $\sigma_{\mathrm{ext}}$($\gamma_{\mathrm{R}}$). MP by comparison shows a less constrained distribution of orientations, with widely distributed tilts $\theta$.
The projected aspect ratios PAR versus the corresponding diameters $D_{\mathrm{p}}$ are also shown in Fig. 5b. The median PAR is $M\{\mathrm{PAR}\}=0.987$ for OL and 0.981 for MP, with its dependence on $g$ given in Fig. 3a. These values are significantly higher than those of MAR. The histograms of $D_{\mathrm{p}}$ are shown and the median for MP is $M\{D_{\mathrm{p}}\}=27.02\,\mathrm{nm}$, larger than $25.37,\mathrm{nm}$ for OL. Remarkably, both PAR and $D_{\mathrm{p}}$ for MP are in good agreement with the TEM results ($M\{\mathrm{PAR}\}=0.985$, $M\{D_{\mathrm{p}}\}=27.75$ nm).
Furthermore, for the two rounder particles in Fig. 5c, OL shows distinctly bimodal distributions , while MP returns unimodal distributions. For these examples, at one standard deviation from the mean, the precision in $D_{V}$ is improved to $(0.04-0.07)\,\mathrm{nm}$, and the precision in AR is (0.37% - 2.8%) for OL and (0.45%-2%) for MP, depending on the NP.
@@ V Conclusions
We have presented an enhanced ON method with optical cross-sections measured using six spectral channels, and both linear and radial polarisation to probe both in-plane and out-of-plane NP response. For nominally $30\,\mathrm{nm}$ ultra-uniform gold nanospheres, we demonstrated significant improvement in the precision and accuracy of the method. Precision in the size and shape has been improved to below $0.1\,\mathrm{nm}$ and $5\%$, respectively. We examined the effect of the permittivity dataset on the morphometry retrieval and found that McPeak et al.[46] with surface damping factor $g=1.8$ yields the best results. We find that the small effect of retardation for these NP, which are some 12 times smaller than the light wavelength in the surrounding medium, provides a LSPR red-shift of some 5 nm. This is significant due to the high precision of the measurements, and correcting for it improves accuracy to within 5% in shape and size using TEM measurements as reference. We however note that this accuracy is contingent on the permittivity used to describe the material response.
Future development could see the use of numerical models, capable of treating particles outside of the limitations of analytical models. Another approximation limiting the accuracy is the use of a local permittivity to describe the nanoparticle optical response, which is known to break down on the sub-nm scale in general. Again, more complex theory models[56] can be used to improve on this. The precision of the method is so high that small systematic effects in the experimental setup, such as the angular resolved transmission and the angular intensity distributions, are relevant to achieve an accuracy comparable to, or better than, the precision. These can be corrected by further characterisation of the instrumentation.
Overall, this study shows the ability of the ON to complement TEM, and as a stand-alone methodology for accurate NP morphometric studies. It also offers an interesting approach to determine the permittivity of materials on the nanoscale, and to uncover non-local effects.